OpenStudy (anonymous):

completing the square x ^2 +6x -8 =0

OpenStudy (johnweldon1993):

You would first add 8 to both sides to get it in the correct form \[\large x^2 + 6x = 8\] You now have an equation in the form \[\large ax^2 + bx = c\] You take your 'b' number (6) divide it by 2...and square the result...and add that to both sides of the equation..what do you get when you do that?

OpenStudy (anonymous):

you get 9 and add to bith sides ?

OpenStudy (johnweldon1993):

Right! you add 9 to both sides...so now we have \[x^2 + 6x + 9 = 8 + 9\] Next step....factor that left hand side of the equation...

OpenStudy (anonymous):

ooohh can you help me with that ..

OpenStudy (johnweldon1993):

No problem..the thing is you'll see that you get a perfect square.. What 2 numbers multiply to get 9...and also add to get 6? 3 and 3 comes to mind \[\large x^2 + 6x + 9 \space \text {can be factored into} \space (x+3)(x+3)\] Now its dumb to write (x + 3) twice...so we write \[(x + 3)^2 = 17\] What would we do next?

OpenStudy (anonymous):

(X+ 3) ^2 - 17 = 0 ?

OpenStudy (anonymous):

is that the answer then ?

OpenStudy (johnweldon1993):

No you wouldnt bring the 17 over...you would square root both sides of the equation \[x + 3 = \pm \sqrt{17}\] Now you would solve the 2 equations

OpenStudy (anonymous):

oh ok

OpenStudy (johnweldon1993):

so \[x + 3 = \sqrt{17}\] and \[x + 3 = -\sqrt{17}\] Solve for each 'x'

OpenStudy (anonymous):

-3-sqrt(17)

OpenStudy (anonymous):

x = sqrt 17 - 3